OpenStudy (anonymous):

HELP!!!! descartes rule f(x)=x^2+3x+5

OpenStudy (anonymous):

@.Sam. @ajprincess @Eyad @RadEn

OpenStudy (perl):

remind me of descartes rule

OpenStudy (anonymous):

It was determined, by the Fundamental Theorem of Algebra, that there are five values of x that make f(x) = x5 + 2x4 – 4x3 – 4x2 – 5x – 6 equal to zero. These values for x are also known as the zeros of the function. Using the Rational Root Theorem and synthetic division, 2, –1 and –3 were found to be three of the five zeros of the function. Notice that these three zeros, which are also known as roots, are the x-intercepts of the graph of this function.

OpenStudy (anonymous):

@perl

OpenStudy (perl):

ok thanks , one moment

OpenStudy (anonymous):

@perl, your avatar is in extremely poor taste.

OpenStudy (perl):

dont like it?

OpenStudy (anonymous):

Not even a little.

OpenStudy (perl):

sammy, what are the directions to this problem?

OpenStudy (anonymous):

using the descartes rule, solve

OpenStudy (anonymous):

if you dont understand this one, im also stuck with this one... find possible ration root f(x)=2x^3+3x^2-x+4

OpenStudy (perl):

ok one sec , i gotta update my pic

OpenStudy (anonymous):

ok

OpenStudy (perl):

the Rational Root Theorem gives you information about how many rational roots there could be. Descartes' Rule of Signs gives you information about how many positive or negative roots there could be (which may or may not be rational). for example, if p(x) = x^2 + 1, the rational root test tells us that the only rational roots (if they exist) will be -1, and 1. applying the rule of signs, we find that there are no positive roots (because there are no sign changes for p(x)), and there are no negative roots (because there are no sign changes for p(-x), which happens to equal p(x) in this case). so as you can see, the two tests can yield different results.

OpenStudy (perl):

f(x) = x5 + 2x4 – 4x3 – 4x2 – 5x – 6

OpenStudy (perl):

you have 1 sign change, so there is exactly only 1 positive real root

OpenStudy (anonymous):

so how do i do my problem? @perl

OpenStudy (perl):

then plug -x f(-x) = (-x)^5 + 2(-x)^4 - 4(-x)^3 - 5(-x) - 6

OpenStudy (perl):

f(-x) = (-x)^5 + 2(-x)^4 - 4(-x)^3 - 5(-x) - 6 = -x^5 + 2x^4 +4x^3 + 5x - 6

OpenStudy (perl):

now there are 2 sign changes, so there are 2 or zero negative roots. since you found 2 negative roots using synthetic division, so we have a total of 3 roots (and 2 non real roots)

OpenStudy (anonymous):

which problem are you solving for??? @perl

OpenStudy (perl):

oh your 2nd example

OpenStudy (anonymous):

and thats just the final answer? @perl

OpenStudy (perl):

f(x) has no positive real roots, but f(-x) = (-x)^2 +3(-x)+5 = x^2 -3x + 5 has either two negative roots or 0 negative roots

OpenStudy (anonymous):

what about the original problem i posted? @perl

OpenStudy (perl):

well descartes rule says there are either 0 negative real roots or 2 negative real roots

OpenStudy (perl):

and i just checked usin quadratic formula, there are no real roots

OpenStudy (perl):

descartes sign rule says there are no positive real roots, and there are either two or zero negative real roots